A ball of mass 2.00 kg is tied to a string and swung in a horizontal circle of radius 5 m at a speed of 10 m/s. The magnitude of the centripetal force required for its movement is
1
Expert's answer
2019-09-09T11:01:35-0400

The magnitude of the centripetal force


"F=m\frac{v^2}{R}"

"=2.0\:\rm kg\times \frac{(10\: m/s)^2}{5.0\: m}=40\:N"